OpenStudy (samara8954):

Colleen has an ice chest with 20 cans of soda in it. Eight are cola, five are lemon-lime, and the rest are grape. If she randomly pulls a can of soda out of the ice chest, what are the chances that she gets a lemon-lime soda? 25% 35% 40% 80%

OpenStudy (bmk614):

\[\frac{ lemonlime }{ total }=\frac{ 5 }{ 20 }=\frac{ 1 }{ 4 }=.25\]

OpenStudy (samara8954):

can you help me with more?

OpenStudy (bmk614):

Yeah sure! Just tag me them!